Sharkattack123:

HELP Factor Completely. x³ + 9x² + 3x + 27

Sharkattack123:

Is the answer (x + 9) x (x² + 3)

Sharkattack123:

@shoaibchaudhry

Darf87:

the answer would be what you wrote above but no x in the middle hope this heps

Sharkattack123:

@darf87 wrote:
the answer would be what you wrote above but no x in the middle hope this heps
No "x" ?

kacchanforlife:

@sharkattack123 wrote:
Is the answer (x + 9) x (x² + 3)
thats what i got

Darf87:

@sharkattack123 wrote:
@darf87 wrote:
the answer would be what you wrote above but no x in the middle hope this heps
No "x" ?
The answer is (x+9)(x2+3) not (x+9)x(x2+3)

Sharkattack123:

@kacchanforlife wrote:
@sharkattack123 wrote:
Is the answer (x + 9) x (x² + 3)
thats what i got
Ye, just making sure i'm right though.😢

kacchanforlife:

oh i didnt read that right sorry

Sharkattack123:

@darf87 wrote:
@sharkattack123 wrote:
@darf87 wrote:
the answer would be what you wrote above but no x in the middle hope this heps
No "x" ?
The answer is (x+9)(x2+3) not (x+9)x(x2+3)
Alright. How did you get it though.🤔

Darf87:

@sharkattack123 wrote:
@darf87 wrote:
@sharkattack123 wrote:
@darf87 wrote:
the answer would be what you wrote above but no x in the middle hope this heps
No "x" ?
The answer is (x+9)(x2+3) not (x+9)x(x2+3)
Alright. How did you get it though.🤔
this gonna take a while but ill type it out

Sharkattack123:

@darf87 wrote:
@sharkattack123 wrote:
@darf87 wrote:
@sharkattack123 wrote:
@darf87 wrote:
the answer would be what you wrote above but no x in the middle hope this heps
No "x" ?
The answer is (x+9)(x2+3) not (x+9)x(x2+3)
Alright. How did you get it though.🤔
this gonna take a while but ill type it out
Thank you

ShoaibChaudhry:

I got (x+9)(x^2+3)

kacchanforlife:

its how u got it but when ur using parenthese, u dont put the multiplication sign in ur problem

ShoaibChaudhry:

I think Kachaan but X meaning times

kacchanforlife:

yes i know

Sharkattack123:

@sxdsouls wrote:
ur answer is right kasy unless ur trying to learn it
I'm trying to make sure it's right because my teacher is going to ask my questions about it, so I wanna make sure I got the correct answer. </3

ANoob:

Factor by grouping x³ + 9x² + 3x + 27 x^2(x+9)+3(x+9) (x^2+3)(x+9)

Extrinix:

He would indeed be right on the answer, but here's how I was told to do it: Given equation: \(x^3+9x^2+3x+27\) Separate the equation into two: \((x^3+9x^2)+(3x+27)\) Factor each of these out: \(x^2(x+9)+3(x+9)\) Look for commons and cancel them out \(x^2(x+9)+3\) Combine the factored out values: \((x^2+3)(x+9)\)

kacchanforlife:

ur right but just dont put the times sign

Darf87:

@kacchanforlife wrote:
ur right but just dont put the times sign
mhm thats what I meant

Sharkattack123:

Alright thanks

kacchanforlife:

np!

ShoaibChaudhry:

Smh Shark so dum fr fr :>

Sharkattack123:

@shoaibchaudhry wrote:
Smh Shark so dum fr fr :>
Shut </3

ShoaibChaudhry:

@sharkattack123 wrote:
@shoaibchaudhry wrote:
Smh Shark so dum fr fr :>
Shut </3
Sorry bestie Anyways

Extrinix:

You can test your answers by running a 1 as x through the problem \((1)^3+9(1)^2+3(1)+27 = 40\) \(((1)^2+3)((1)+9) = 40\)

Sharkattack123:

@extrinix wrote:
You can test your answers by running a 1 as x through the problem \((1)^3+9(1)^2+3(1)+27 = 40\) \(((1)^2+3)((1)+9) = 40\)
Ah alright, thanks